home printable Archive hribf Physics Contact

RA2. Accelerator System Status

ORIC Operations and Development (B. A. Tatum)

This has been a very difficult year for ORIC with limited operation. Following some difficulties with the extraction and rf systems, operation with proton beam resumed in February. However, this was short-lived. In late February, trimming coil eight (T8) completely failed due to an un-repairable internal water leak.

ORIC was designed with a set of ten concentric water-cooled trimming coils that nominally operate with currents up to 800 Adc. All ten copper coils are contained in a single stainless steel enclosure (see Figure 1). Matching coil assemblies are mounted on each of the two main magnet pole faces. This is not the first trimming coil failure. In fact, there have been several. Trimming coil T10 failed in 1986, 1/3 of T8 failed in 1988 (the coils in each assembly have either 2 or 3 coolant loops), and T9 failed in 1997 prompting us to fabricate a spare set of coils. With each of the prior failures, we were able to determine an alternate tuning configuration that allowed operation to continue. So we prepared for the possibility of another failure, not knowing when it might occur or whether it would prevent ORIC from delivering beam.

When T8 failed this year, there were also concurrent failures of the deflector septum and rf system. After drying the coil assemblies and repairing the extraction and rf systems we attempted to extract beam without T8, which we did, but not with high enough beam energy or intensity. Thus it became necessary to install the spare coils. Replacement of the trimming coils requires about four months, and we are concurrently implementing several other upgrades and performing considerable preventive maintenance. Note that in replacing the trimming coils, all ten coils are replaced at the same time. Thus ORIC will be restored to its original coil configuration, which should lead to improvements in beam extraction efficiency.

The trimming coil replacement project is scheduled for completion in early November. Once ORIC is again operational, beam will first be delivered to IRIS2 to complete commissioning activities with radioactive ion beam.

Figure RA2-1: New ORIC Trimming Coils During Fabrication

RIB Injector Operations and Development (P.E. Mueller)

During the period from 1 January 2010 to 30 June 2010, the 25 MV Tandem Electrostatic Accelerator delivered beams of

  • 53 kpps [23.05 MV 14+/26+ terminal foil / high energy foil stripped] 520-MeV 95% 132Sn,
  • 73 kpps [23.57 MV 15+/26+ terminal foil / high energy foil stripped] 540-MeV 95% 132Sn,
  • 85 kpps [23.28 MV 15+/27+ terminal foil / high energy foil stripped] 550-MeV 95% 132Sn,
  • 90 kpps [23.7 MV 15+/27+ terminal foil / high energy foil stripped] 560-MeV 95% 132Sn,
  • 53 kpps [23.94 MV 15+/28+ terminal foil / high energy foil stripped] 580-MeV 95% 132Sn, and
  • 23 kpps [24.08 MV 15+/29+ terminal foil / high energy foil stripped] 600-MeV 95% 132Sn to the time-of-flight endstation in Beam Line 23, and
  • 800 kpps [20.93 MV 5+ terminal gas & foil stripped] 117-MeV 26gAl to Beam Line 41.

    The high purity tin beams were produced via proton induced fission of 238U by bombarding a pressed powder uranium carbide target coupled to an Electron Beam Plasma (positive) Ion Source (EBPIS) with 9-10 uA of 50-MeV 1H and passing a positive tin sulfide beam through the recirculating cesium jet charge exchange cell and selecting the negative tin beam resulting from molecular breakup. The 26gAl beam was produced with the multisample cesium sputter negative ion source using pressed powder copper targets loaded with 26gAl2O3.

  • Tandem Operations and Development (M. Meigs)

    The Tandem Accelerator was operated for more than 2100 hours since the last report.   The machine ran at terminal potentials of 2.85 to 24.34 MV.  IRIS1 provided 164 hours of the radioactive beams 26Al and 132Sn.  While the UC source was still on IRIS1, two shifts of 238U were provided.  In addition, the SNICS was used to provide the stable beams 1H, 16O, 19F, 27Al, 32S, 54Fe, 58Ni, 74,76Ge, 76Se, 124Sn and 122,130Te.  About 88 hours were spent conditioning to assure operation above 24 MV.  No tank openings were required during this period.

    Home printable Archive hribf Physics ornl contact